OpenStudy (help_people):

Two sets of equatic expressions are shown below in various forms: Line 1: x2 + 3x + 2 (x + 1)(x + 2) (x + 1.5)2 − 0.25 Line 2: x2 + 5x + 6 (x + 2)(x + 3) (x + 2.5)2 + 6.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (help_people):

@DanJS

OpenStudy (danjs):

keep the first entries the same, expand out the second and third and compare those to the first

OpenStudy (help_people):

how would or do that?

OpenStudy (danjs):

(x + 1)(x+ 2) = x^2 + 2x + x + 2 = x^2 + 3x + 2 that one matches

OpenStudy (danjs):

(x+1.5)^2 - 0.25 = x^2 + 1.5x + 1.5x + 2.25 - 0.25 = x^2 + 3x + 2 That one is also the same, so line 1 is all same

OpenStudy (danjs):

test the second line

OpenStudy (help_people):

ok, but I'm still not sure how to test the second line but i will give it a shot

OpenStudy (help_people):

yea I'm not sure the 1st line looks easier this one does not trust me I'm not trying for you to do my work but I'm still not sure how to do the second lien the first line looked a little easier though :/ @DanJS

OpenStudy (danjs):

\[(x+2.5)^2 + 6.5 = x^2+2.5x+2.5x+6.25 + 6.25\]

OpenStudy (danjs):

that doesn't simplify to the same as the first entry line 2

OpenStudy (help_people):

??? @DanJS

OpenStudy (help_people):

are you saying the answer a ?

OpenStudy (help_people):

@welshfella

OpenStudy (danjs):

yes

OpenStudy (help_people):

wait so line b I'm sorry just confirming @DanJS

OpenStudy (help_people):

i mean to say lien 1 @DanJS

OpenStudy (danjs):

yea